OpenStudy (mathisfun13):

Find an irrational numbers between 4/9 and 5/9.

OpenStudy (photon336):

I guess we would have to find a decimal with a number that doesn't repeat, or terminate.

OpenStudy (daniel.ohearn1):

Is it find a irrational number or all irrational numbers between those fractions?

jimthompson5910 (jim_thompson5910):

It's impossible to list out all of the irrational numbers between any two real numbers

OpenStudy (mathisfun13):

.4510100100010000.. is one

OpenStudy (mathisfun13):

because it is greater than .4444... and less than .5555...

jimthompson5910 (jim_thompson5910):

correct @mathisfun13

OpenStudy (mathisfun13):

I just need another example

jimthompson5910 (jim_thompson5910):

another one is step 1) take the midpoint of 0.4444... and 0.5555.... midpoint formula is M = (a+b)/2 step 2) add on some very small irrational number, like pi/1000 to make the overall result irrational

jimthompson5910 (jim_thompson5910):

or you can compute the geometric mean of 4/9 and 5/9

OpenStudy (across):

gee... this cannot be easier...\[\frac{16}{81}<\frac{21}{81}<\frac{25}{81}\\\frac{4}{9}<\frac{\sqrt{21}}{9}<\frac{5}{9}\]

jimthompson5910 (jim_thompson5910):

that works too